OpenStudy (anonymous):

Suppose that to make the golf team you need to score no more than 74 on average over 5 games. If you scored 77, 71, 77, and 67 in your first 4 games what is the highest score you can shoot in your 5th game and still make the team?

OpenStudy (callisto):

let x be the highest score. 77+71+77+67 +x --------------- = 74 5 292 +x = 74(5) Can you work from here?